The PTO is selling raffle tickets to raise money for classroom supplies. A raffle ticket costs $5. There is 1 winning ticket out of the 290 tickets sold. The winner gets a prize worth $72. Round your answers to the nearest cent.
What is the expected value (to you) of one raffle ticket? $

Calculate the expected value (to you) if you purchase 6 raffle tickets. $

What is the expected value (to the PTO) of one raffle ticket? $

If the PTO sells all 290 raffle tickets, how much money can they expect to raise for the classroom supplies? $

Answer by Boreal(15235)

The expected value is $72*1/290-$5 for the ticket
This is $0.248-$5 or -$4.75 expected value
-
The expected value of 1 raffle ticket to the PTO is +$4.75
-
NOT to win is 289/290*288/289*287/288*286/287*285/286*284/285=0.9793
So probability of winning is complement or 0.0207
Expected value is 0.0207*$72-S5=-$3.51
-
They will take in 290*5=$1450-$72 paid out or $1378.